A hydrogen cylinder is designed to withstand an internal pressure of 100atm .at 27 degree c ,hydrogen is pumped into cylinder which exerts a pressure of 20atm .at what temperature does the danger explosion first sets in ???????????????????????
1
Expert's answer
2013-11-26T05:09:22-0500
if we suppose that process is isochoric, then

p/T = const, p1/T1 = p2/T2

p1= 20 atm, p2= 100 atm(maximal non-dangerous pressure), T1= 27 C = 300 K. Hence

T2 = p2/p1 * T1 = 100/20*300 = 5*300 = 1500 K = 1227 C

answer is at 1227 C degrees.